differentiate implicity the following equations. show complete solutions



1. e^(2x+3y) =x^2 - In(xy^3)



2. x^2 y^3 = xy+x^3 + y^2 + 5

1
Expert's answer
2021-12-27T16:18:19-0500

"1.\ e^{(2x+3y)} =x^2 - In(xy^3)"

Differentiating both sides w.r.t "x"

"e^{(2x+3y)}\{2+3y'\} =2x-\frac{1}{xy^3}.[y^3+2xy^2y']"


"2. x^2 y^3 = xy+x^3 + y^2 + 5"

Differentiating both sides w.r.t "x"

"2xy^3+2x^2y^2y'=y+xy'+3x^2+2yy'"